OpenStudy (anonymous):

Two consecutive sides of a parallelogram measure 32 centimeters and 15 centimeters. The angle made by the pair of consecutive sides is 135°. What is the area of the parallelogram?

OpenStudy (anonymous):

draw the height from a vertex of the parallelogram on the 32 cm base

OpenStudy (anonymous):

94 is the area

OpenStudy (anonymous):

16.6 square centimeters 339.4 square centimeters 169.7 square centimeters 678.8 square centimeters

OpenStudy (anonymous):

sigh the answer is 339.4

OpenStudy (anonymous):

to find the height (h) of the parallelogram, find: sin 45 = h/15 h = (15√2)/2 cm so, Area of parallelogram = (32) * (15√2)/2 = 240√2 cm^2 ≈ 339.411 cm^2

OpenStudy (anonymous):

@AleshGames is wright I forgot a step